OpenStudy (anonymous):

Four fair dice are rolled at the same time. What is the probability that at least one of the dice will show 3?

OpenStudy (anonymous):

probability of showing 3 on one die=1/6 Probability of not showing 3 on one die=5/6 Probability of not showing 3 on any die=5/6*5/6*5/6*5/6=625/1296 Probability of showing 3 at least once=1-625/1296=671/1296